# Efficient Manipulation-Enhanced Semantic Mapping With Uncertainty-Informed Action Selection

### v1

@@ III-C Push Sampling
To enhance visibility and maximize information gain in cluttered and confined environments, particularly when only executing observations proves insufficient, we leverage strategic uncertainty-aware object manipulation to uncover occluded regions. In the following, we detail the process of identifying suitable manipulation targets, estimating visibility corridors, determining effective push directions, and finally sampling feasible push interactions. An illustration of each step is shown in Fig. 3.
@@ III-C1 Target Locations
Unlike existing object-search methods [23, 3], we define target locations as regions of the map with high semantic uncertainty. We prioritize semantic uncertainty over occupancy uncertainty because even when occupancy estimates are confident, class labels can remain ambiguous. Focusing on semantics drives the robot to discover new objects and improve its understanding of known ones. To identify these regions, we compute a distance transform over a thresholded version of the semantic uncertainty, where cells with an semantic uncertainty below $0.1$ are considered sufficiently certain and set to 0. The distance transform assigns each cell a score based on how far it is from the nearest confidently classified cell (either free or occupied), with the score scaled by cells occupancy uncertainty value. From the resulting distance map, the a maximum offive cells with the highest distance values are selected as candidate target locations, while ensuring a minimum pairwise distance to promote spatial diversity as highlighted in Fig. 3b.
@@
Figure 3: Uncertainty-informed push sampling. From the current map belief (a), with shelf borders indicated in brown, we generate target points from a distance map representing the occupancy uncertainty-weighted distance to the nearest confidently classified cell (b). Then, we compute visibility corridors using raycasting from each target location to the front of the shelf, identifying occluding objects, as highlighted in yellow for one of the corridors (c). Finally, $n_{p}$ push proposals are sampled from feasible start regions along the identified occluding object, aligned with pushing corridors, that represent the best direction for the object to be pushed into (d). The generated push samples are used as input to the manipulation CNABU to evaluate their expected information gain.
@@ III-C2 Visibility Corridor
For each selected target location, a visibility corridor is constructed to identify potential occlusions. This is accomplished using raycasting, from the target point 180 degrees toward the front of the shelf, recording the number of semantic object instances from the one-hot-encoded map that are encountered along the ray as well as the average occupancy probability weighted by the epistemic occupancy uncertainty. Neighboring rays that exhibit similar spatial structures (in terms of object instances and non-occupied space) are clustered into unified corridors up to a maximum radius of $n_{c}$ cells. These corridors are scored according to four criteria:
$score=k1\cdot w+k2\cdot l+k3\cdot\bar{p}_{\text{occ}}+k4\cdot n_{\text{occ\_obj}}$ (4)
where $w$ is the width of the corridor, $l$ the length of the visibility corridor, $\bar{p}_{\text{occ}}$ the average occupancy probability, $n_{\text{occ\_obj}}$ the number of occluding objects, and $k_{1,2,3,4}$ are weights. The object closest to the front of the shelf within the corridor with the highest score is selected as the manipulation target (see Fig. 3c). Note that if no occluding object is found, the target is considered observable through future viewpoint planning and no push is considered.
@@ III-C3 Pushing Corridor
The objective is to find a direction that minimizes secondary pushing of adjacent objects and prefers already mapped free areas over highly uncertain ones. Therefore, unlike the visibility corridor, sectors with lower average occupancy and greater clearance are preferred to ensure good reachability and low collision probabilities with other objects. To determine viable pushing directions for the selected object, again raycasting is employed. Rays are radially cast from the centroid of the object identified as pushing target toward a circular boundary with a maximum radius of $n_{c}$ cells. We identify object entities and their centroids by clustering connected regions of similar semantic labels in the current map. For each object, we cast rays from its centroid inside the shelf in multiple directions, grouping these rays into sectors with an maximum angle of 30 degrees. Each sector is evaluated based on the average occupancy along its rays and spatial constraints such as nearby obstacles or shelf boundaries (see Fig. 3d, blue corridor).
III-C4 Sampling Push Interactions
To generate possible pushes to execute, a starting corridor is established on the side opposite the pushing corridor (see Fig. 3d). This is achieved by identifying a wider region on the object’s contour based on the intersection points of lines defined by the endpoints of both corridors.
@@ III-D Manipulation-Enhanced Planning
We combine our viewpoint selection policy with the uncertainty informed push-based manipulation strategy to realize a complete manipulation-enhanced mapping pipeline. While the RL-based next-best view (NBV) planning agent seeks to maximize information gain through passive camera movement, physical occlusions can make certain regions permanently inaccessible to observation. In such cases, the system queries the push sampling module to actively remove occlusions.
For decision making, we employ a variant of the occlusion-aware Volumetric Information Gain (VIG) planning strategy proposed by [10]. For the NBV module, we query the trained RL agent to generate a single candidate action, obtain the predicted camera pose, and compute the associated VIG based on the evidential occupancy prediction at time $t-1$. Similarly, for the push sampling module, we simulate the outcome of each push by generating its parameterization and input it together with the evidential occupancy prediction at time $t-1$ to the manipulation-CNABU to predict the resulting updated evidential map. Subsequently, we query the NBV agent on the post-push map, incorporating the effects of the simulated manipulation, and compute the corresponding VIG.
The system then compares the maximum expected VIG from viewpoint planning ($\text{VIG}_{nbv}$) against that from push actions ($\text{VIG}_{push}$). If the VIG achievable through viewpoint selection exceeds that of any proposed push by a predefined margin $\text{VIG}_{nbv}\cdot\Delta_{\text{view}}>\text{VIG}_{push}$, the system executes the corresponding NBV action. Otherwise, if pushing is expected to yield higher information gain, the robot proceeds to execute the selected push to actively modify the scene and takes its predicted belief generated by the manipulation-CNABU as map representation. [TODO: does an observation action follow or not? This is also unclear from Fig 2]
This strategy ensures that executing viewpoints is always prioritized when sufficient information gain can be achieved without manipulation, while empowering the system to physically interact with the environment whenever necessary to uncover previously inaccessible regions and maximize overall map accuracy.
@@ IV Experimental Evaluation
We perform three experimental evaluations to showcase the performance of our next-best view and push selection approach. First, we present quantitative simulation results that highlight our pipeline’s improvements in mapping accuracy, compared to the state-of-the-art for manipulation-enhanced mapping [10]. Next, we present several ablations of our method to highlight the influence of uncertainty measurements for action selection and evaluate different versions of our novel push sampling strategy. Finally, we show a qualitative experiment in a real hardware setup in zero-shot fashion in the supplementary video.
@@
[…]
RL-CNABU-View (Ours) Active None $0.844\pm 0.001$ $0.607\pm 0.002$
RL-CNABU-View-Privileged Active None $0.846\pm 0.001$ $0.609\pm 0.002$
POMDP-CNABU-View [10] Active None $0.862\pm 0.001$ $0.641\pm 0.001$
Random-View Active None $0.839\pm 0.002$ $0.595\pm 0.001$
POMDP-CNABU-Push [10] Interactive Random Sampling $0.877\pm 0.003$ $0.720\pm 0.019$ $1.927\pm 0.046$ $5.25\pm 2.61$ 25.33%
POMDP-CNABU-Push (DC) Interactive Random Sampling $0.832\pm 0.005$ $0.638\pm 0.022$ $1.841\pm 0.033$ $4.65\pm 2.50$ 25.33%
TABLE II: Quantitative comparison over 25 cluttered scenes, averaged across three trials. We report occupancy and semantic mIoU, mean action displacement (mAD), push count, and collision rate. Our informed strategies ($u_{s}|u_{o}$, $u_{s}|u_{b}$, $u_{o}|u_{o}$) and RL-based methods show benefits of uncertainty-aware action selection. While POMDP-CNABU-Push achieves the highest semantic mIoU, it does so at the cost of more frequent pushing and a high collision rate. In contrast, our RL-CNABU-Push approach offers a better balance, achieving competitive mapping performance with fewer pushes, lower movement, and reduced number of collisions.
@@ IV-A Experimental Setup and Training
[…]
The following hyperparameters are used for all experiments: $N_{\mathrm{hist}}=4$, $\Delta_{\text{view}}=2$, $\theta_{\text{sem}}=0.65$, and $\theta_{\text{occ}}=0.87$. For the reward function of the viewpoint planning agent, we set $\alpha=10$ and $\beta=2$.
@@ IV-B Metrics and Baselines
  • mIOU: Semantic and occupancy mean Intersection-over-Union.
  • mAD: Mean displacement of each object averaged over all evaluated scenes.
@@
  • Timings: Inference and execution timings of the pipelines
@@
  • POMDP-CNABU-Push (DC): POMDP-CNABU-Push terminated when any of the objects drop down.
@@ IV-C Quantitative Experiments
We present our quantitative experiments, designed to evaluate the effectiveness of our proposed manipulation-enhanced mapping pipeline. To compare our approach against POMDP-CNABU planners, we re-implemented their pipeline and used the provided CNABU-models from their code release. In order to use the models for our approach, no fine-tuning on the novel set of continuous viewpoints was performed. However, we increased the certainty threshold for full map completion after which no pushes are performed from 95%, as stated in [10], to 99%. This increases the chance of further pushes in the long run and shows the effect of successful push prediction on the over- or under-confidence of the CNABU-models.
IV-C1 Timings
We evaluate the computational efficiency of our methods against the POMDP-CNABU baselines. As shown in Table III, our method significantly reduces decision latency in both viewpoint and manipulation selection. This is primarily due to our reinforcement learning-based viewpoint planner, which operates in a continuous action space and avoids exhaustive enumeration and ray-casting of discrete candidates, as well as our informed push sampling strategy, which narrows down the set of evaluated push actions while preserving their impact on map improvement. In particular, our RL-CNABU-View planner achieves a 99.7% reduction in average inference time per step compared to the original POMDP-CNABU-View planner. Similarly, the informed push selection time and the total per-iteration execution times are significantly reduced, enabling faster scene coverage and improved responsiveness.
IV-C2 Map Quality
To assess the effectiveness of our proposed system in building complete and accurate scene representations, we evaluate semantic and occupancy map quality across 25 highly cluttered shelf environments with an action budget of 40 steps and evaluate them according to the metrics specified in Sec. IV-B. The results across the whole action budget is illustrated in Fig. 4 with a more detailed analysis of the last step shown in Tab. II. In the following, we summarize the main results.
Among active perception methods, although POMDP-CNABU-View performs better than RL-CNABU-View in mapping accuracy, it highly relies on the quality of the predefined fixed views and is computationally expensive as it uses a greedy search over up to 300 fine-tuned views, querying expected information gain at each step. In contrast, our RL-CNABU-View predicts the next-best view with a single forward pass through a learned continuous-space policy, substantially reducing computation time while maintaining comparable mIoU.
With regards to active versus interactive perception, Tab. II shows that the worst performing interactive perception method achieves higher semantic mIOU than the best active perception method, highlighting the need for physical interaction. In particular, our uncertainty-informed push method, RL-CNABU-Push achieves higher semantic mIoU than both RL-CNABU-View and POMDP-CNABU-View, demonstrating that targeted push actions can improve mapping accuracy by uncovering hidden regions. For interactive perception, Fig. 4 and Tab. II show that POMDP-CNABU-Push achieves higher mIOU metrics compared to RL-CNABU-Push. However, unlike all other methods, this proposed pipeline [10] does not consider any early stopping when an object drops, e.g., out of the shelf, due to collisions. As can be seen, when terminating after an object drop occured, our proposed RL-CNABU-Push achieves better occupancy and semantic mIOU compared to POMDP-CNABU-Push (DC).
As all POMDP-CNABU-Push variants use random push sampling to generate pushes without knowledge of effective pushing corridors, they result in the highest mean displacement and collision rates among interactive perception methods. In POMDP-CNABU-Informed, replacing random sampling with our uncertainty-informed strategy increases mIoU while significantly reducing average displacement, number of pushes, and collisions. The reduced displacement and collisions in all the informed push methods confirm that our interactions are minimally invasive compared to the random push sampling baseline. Furthermore, informed push sampling methods require only 2–3 pushes versus 4–5 in random methods to achieve comparable or better map quality, demonstrating more targeted uncovering of hidden spaces.
These results confirm that leveraging epistemic uncertainty for both viewpoint and manipulation planning leads to comparable or better quality map quality, particularly in scenes with high occlusion. Furthermore, as the timing results indicate our method achieves comparable mapping accuracy and reliability, with significantly better computational efficiency
@@ IV-C2 Map Quality
TABLE III: Timing performance in seconds across the 25 evaluation scenes averaged over 40-step rollouts and across three trials. Our RL and uncertainty aware push approach highly reduces the individual and overall inference time of the task.

### v2

@@ III-C Push Sampling
To enhance visibility and maximize information gain in cluttered and confined environments, particularly when only active sensing proves insufficient, we leverage strategic uncertainty-aware object manipulation to uncover occluded regions. In the following, we detail the process of identifying suitable manipulation targets, estimating visibility corridors, determining effective push directions, and finally sampling feasible push interactions. An illustration of each step is shown in Fig. 3.
@@ III-C1 Target Locations
Unlike existing object-search methods [23, 3], we define target locations as regions of the map with high semantic uncertainty. We prioritize semantic uncertainty over occupancy uncertainty because even when occupancy estimates are confident, class labels can remain ambiguous. Focusing on semantics drives the robot to discover new objects and improve its understanding of known ones. To identify these regions, we compute a distance transform over a thresholded version of the semantic uncertainty, where cells with a semantic uncertainty below $0.1$ are considered sufficiently certain and set to 0. This value is empirically chosen to be optimal in terms of isolating regions whose certainty cannot be substantially improved given inherent network imperfections or sensor noise. The distance transform assigns each cell a score based on how far it is from the nearest confidently classified cell (either free or occupied), with the score scaled by cells occupancy uncertainty value. From the resulting distance map, the a maximum of five cells with the highest distance values are selected as candidate target locations, while ensuring a minimum pairwise distance to promote spatial diversity as highlighted in Fig. 3b.
@@
Figure 3: Uncertainty-informed push sampling. From the belief map (a; shelf borders in brown), we build an uncertainty-weighted distance field and sample target points (b). Raycasting from each target to the shelf front yields visibility corridors and associated occluding objects (c; yellow). For each occluder, we generate $n_{p}$ feasible push proposals aligned with the corridor toward the clearing direction (d). Afterwards, the manipulation CNABU evaluates the expected information gain of the push proposals.
@@ III-C2 Visibility Corridor
For each selected target location, a visibility corridor is constructed to identify potential occlusions. This is achieved via raycasting from the target point towards the front of the shelf, defined by a fixed $y$ value and a variable $x$ value in the range [0, maximum width]. Along each ray, the number of semantic object instances from the one-hot-encoded map encountered is recorded, as well as the average occupancy probability, weighted by the corresponding epistemic occupancy uncertainty.
Neighboring rays that exhibit similar spatial structures (in terms of the order of object instances and non-occupied space) are clustered into unified corridors up to a maximum radius of $n_{c}$ cells. Each corridor is represented as a structure of its start point, width, length, order of objects, and average occupancy probability. These corridors are scored according to four criteria:
$score=k_{1}\cdot w+k_{2}\cdot l+k_{3}\cdot\bar{p}_{\text{occ}}+k_{4}\cdot n_{\text{occ\_obj}}$ (5)
where $w$ is the width of the corridor, $l$ the normalized length of the visibility corridor, $\bar{p}_{\text{occ}}$ the average occupancy probability scaled by the normalized amount of unknown cells, $n_{\text{occ\_obj}}$ the number of occluding objects, and $k_{1,2,3,4}$ are weights empirically found. The object closest to the front of the shelf within the corridor with the highest score is selected as the manipulation target (see Fig. 3c). Note that if no occluding object is found, the target is considered observable through pure future viewpoint planning.
@@ III-C3 Pushing Corridor
The objective is to find a direction that minimizes secondary or inadvertent pushing of adjacent objects and prefers already mapped free areas over highly uncertain ones. Therefore, unlike the visibility corridor, sectors with lower average occupancy and greater clearance are preferred to ensure good reachability and low collision probabilities with other objects. To determine viable pushing directions for the selected object, again raycasting is employed. Rays are radially cast from the centroid of the object identified as pushing target toward a circular boundary with a maximum radius of $n_{c}$ cells. We identify object entities and their centroids by clustering connected regions of similar semantic labels in the current map. For each object, we cast rays from its centroid inside the shelf in multiple directions, grouping these rays into sectors with an maximum angle of $30^{\circ}$. Each sector is evaluated based on the average occupancy along its rays and spatial constraints such as nearby obstacles or shelf boundaries (see Fig. 3d, blue corridor).
III-C4 Sampling Push Candidates
To generate possible pushes to execute, a starting corridor is established on the side opposite the pushing corridor (see Fig. 3d). This is achieved by extending the lines from the end points of the corridor through the crossed contour intersection points, towards the opposite side of the contour.
@@ III-D Manipulation-Enhanced Planning
We combine our viewpoint selection policy with the uncertainty informed push-based manipulation strategy to realize a complete manipulation-enhanced mapping pipeline. While the RL-based next-best view planning agent seeks to maximize information gain through passive camera movement, physical occlusions can make certain regions permanently inaccessible to observation. In such cases, the system queries the push sampling module to actively remove occlusions.
For decision making, we employ a variant of the occlusion-aware Volumetric Information Gain (VIG) method proposed by [10]. For the NBV module, we query the trained RL agent to generate a single candidate action, obtain the predicted camera pose, and compute the associated $\text{VIG}_{nbv}$ based on the evidential occupancy prediction at time t1t-1. Similarly, for the push sampling module, we simulate the outcome of each push by generating its parameterization as described in [10] and input it together with the evidential occupancy prediction at time t1t-1 to the manipulation-CNABU to predict the resulting updated evidential map. Subsequently, we query the NBV agent on the post-push map, incorporating the effects of the simulated manipulation, and compute the corresponding $\text{VIG}_{push}$.
The system then compares the maximum expected $\text{VIG}_{nbv}$ from viewpoint planning against $\text{VIG}_{push}$. If $\text{VIG}_{nbv}$ exceeds that of any proposed push by a predefined margin $\text{VIG}_{nbv}\cdot\Delta_{\text{view}}>\text{VIG}_{push}$, the system executes the corresponding NBV action. Otherwise, if pushing is expected to yield higher information gain, the robot proceeds to execute the selected push to actively modify the scene and takes its predicted belief generated by the manipulation-CNABU as map representation. This strategy ensures that executing viewpoints is always prioritized when sufficient information gain can be achieved without manipulation, while empowering the system to physically interact with the environment whenever necessary to uncover previously inaccessible regions and maximize overall map accuracy.
Parameter Value Parameter Value
$N_{\mathrm{hist}}$ 4 $\Delta_{\text{view}}$ 2
$\theta_{\text{sem}}$ 0.65 $\theta_{\text{occ}}$ 0.87
$\gamma_{p}$ 0.5 $\gamma_{e}$ 0.5
$\theta_{p}$ 0.1 m $\theta_{r}$ 0.034 rad
$w_{1}$ 10 $w_{2}$ 2
Pushing Parameters
$k_{1}$ 2 $k_{2}$ 3
$k_{3}$ 4 $k_{4}$ 5
TABLE II: Hyperparameters for RL training and push sampling.
@@ IV Experimental Evaluation
We perform three experimental evaluations to showcase the performance of our next-best view and push selection approach. First, we present quantitative simulation results that highlight our pipeline’s improvements in mapping accuracy, compared to the state-of-the-art for manipulation-enhanced mapping [10]. Next, we present several ablations of our method to highlight the influence of uncertainty measurements for action selection and evaluate different versions of our novel push sampling strategy. Finally, we show a qualitative experiment in a real hardware setup in zero-shot fashion in the supplementary video [^2] https://youtu.be/9h2MKLzQl80.
@@ IV-A Experimental Setup and Training
The agent is trained for $500,000$ iterations across approximately $12,500$ procedurally generated shelf configurations. Object counts range from 15 to 30 per scene. To simulate varying occlusion levels for training, large objects are biased toward the front of the shelf and smaller ones toward the back, increasing visual complexity. The action space is sampled from a Cartesian space of $0.8\text{\,}\mathrm{m}$\times$0.2\text{\,}\mathrm{m}$\times$0.2\text{\,}\mathrm{m}$$ outside and inside the shelf, respectively. The hyperparameters used for all experiments are stated in Tab. II.
@@ IV-B Metrics and Baselines
  • Timings: Inference and execution timings of the pipelines
  • mIOU: Semantic and occupancy mean Intersection-over-Union, computed as the average per-class IoU between predictions and ground truth, calculated on the CNABU network output after transformation into the map representation.
  • mAD: Mean displacement of add objects averaged over all evaluated scenes.
@@
  • Collisions: Average percentage of collisions occurred over all evaluated scenes.
@@
  • POMDP-CNABU-Push (ToF): POMDP-CNABU-Push terminated when any of the objects falls over.
@@ IV-C Quantitative Experiments
We present our quantitative experiments, designed to evaluate the effectiveness of our proposed manipulation-enhanced mapping pipeline. To compare our approach against the POMDP-CNABU planners [10], we re-implemented their pipeline and used the provided CNABU-models from their code release. In order to use the models for our approach, no fine-tuning on the novel set of continuous viewpoints was performed. However, we increased the certainty threshold for full map completion after which no pushes are performed from 95%, as stated in [10], to 99%. This increases the chance of further pushes in the long run and shows the effect of successful push prediction on the over- or under-confidence of the CNABU-models.
To evaluate the effectiveness of our system in creating accurate scene representations, we assess the quality of semantic and occupancy maps across 25 highly cluttered shelf environments with an action budget of 40 steps. We evaluate them according to the metrics specified in Sec. IV-B.
@@
TABLE III: Timing performance in seconds across the 25 evaluation scenes averaged over 40-step rollouts and across three trials. Our RL and uncertainty aware push approach highly reduces in general the inference times of the task.
IV-C1 Timings
As shown in Table III, our method significantly reduces decision latency in both viewpoint and manipulation selection. This is primarily due to our reinforcement learning-based viewpoint planner, which operates in a continuous action space and avoids exhaustive enumeration and ray-casting of discrete candidates, as well as our informed push sampling strategy, which narrows down the set of evaluated push actions while preserving their impact on map improvement. In particular, our RL-CNABU-View planner achieves a 99.7% reduction in average inference time per step compared to the original POMDP-CNABU-View planner. Similarly, the informed push selection time and the total per-iteration execution times are significantly reduced, enabling faster scene coverage and improved responsiveness.
Model Perception Push Strategy Occ. mIOU $\uparrow$ Sem. mIOU $\uparrow$ mAD [m] $\downarrow$ NumPush [#] $\downarrow$ Collisions $\%$] $\downarrow$
RL-CNABU-View (Ours) Active None $0.844\pm 0.001$ $0.607\pm 0.002$ - - -
RL-CNABU-View-Privileged Active None $0.846\pm 0.001$ $0.609\pm 0.002$ - - -
POMDP-CNABU-View [10] Active None $0.862\pm 0.001$ $0.641\pm 0.001$ - - -
Random-View Active None $0.839\pm 0.002$ $0.595\pm 0.001$ - - -
POMDP-CNABU-Push [10] Interactive Random Sampling $\bf{0.877\pm 0.003}$ $\bf{0.720\pm 0.019}$ $1.927\pm 0.046$ $5.25\pm 2.61$ 25.33%
POMDP-CNABU-Push (ToF) Interactive Random Sampling $0.832\pm 0.005$ $0.638\pm 0.022$ $1.841\pm 0.033$ $4.65\pm 2.50$ 25.33%
TABLE IV: Quantitative comparison over 25 cluttered scenes, averaged across three trials. We report occupancy and semantic mIoU, mean action displacement (mAD), push count, and collision rate. Our informed strategies ($u_{s}|u_{o}$, $u_{s}|u_{b}$, $u_{o}|u_{o}$) and RL-based methods show benefits of uncertainty-aware action selection. While POMDP-CNABU-Push achieves the highest semantic mIoU, it does so at the cost of more frequent pushing and a high collision rate. In contrast, our RL-CNABU-Push approach offers a better balance, achieving competitive mapping performance with fewer pushes, lower object displacement, and reduced number of collisions.
IV-C2 Map Quality
The results in terms of map quality across the whole action budget of 40 steps is illustrated in Fig. 4 with a more detailed analysis of the last step shown in Tab. IV. In the following, we summarize the main results.
Among active perception methods, although POMDP-CNABU-View performs better than RL-CNABU-View in mapping accuracy, it highly relies on the quality of the predefined fixed views and is computationally expensive as it uses a exhaustive greedy search over up to 300 fine-tuned views, querying expected information gain at each step. In contrast, our RL-CNABU-View predicts the next-best view with a single forward pass through a learned continuous-space policy, substantially reducing computation time, as discussed in Sec. IV-C1, while maintaining comparable mIoU.
With regards to active versus interactive perception, Tab. IV shows that the worst performing interactive perception method achieves higher semantic mIOU than the best active perception method, highlighting the need for physical interaction. In particular, our uncertainty-informed push method, RL-CNABU-Push achieves higher semantic mIoU than both RL-CNABU-View and POMDP-CNABU-View, demonstrating that targeted push actions can improve mapping accuracy by uncovering hidden regions. For interactive perception, Fig. 4 and Tab. IV show that POMDP-CNABU-Push achieves higher mIOU metrics compared to RL-CNABU-Push. However, unlike all other methods, this proposed pipeline [10] does not consider any early stopping when an object falls over, e.g., out of the shelf, due to collisions. When POMDP-CNABU-Push is terminated on an object fall, referred to as POMDP-CNABU-Push (ToF), our proposed RL-CNABU-Push achieves higher occupancy and semantic mIOU compared to POMDP-CNABU-Push (ToF).
As all POMDP-CNABU-Push variants use random push sampling to generate pushes without knowledge of effective pushing corridors, they result in the highest mean displacement and collision rates among interactive perception methods. In POMDP-CNABU-Informed, replacing random sampling with our uncertainty-informed strategy increases mIoU while significantly reducing average displacement, number of pushes, and collisions. The reduced displacement and collisions in all the informed push methods confirm that our interactions are minimally invasive compared to the random push sampling baseline. Furthermore, informed push sampling methods require only 2–3 pushes versus 4–5 in random methods to achieve comparable or better map quality, demonstrating more targeted uncovering of hidden spaces. These results confirm that leveraging uncertainty for both viewpoint and manipulation planning leads to comparable or better quality map quality, particularly in scenes with high occlusion. Furthermore, as the timing results indicate, our method achieves comparable mapping accuracy, with significantly reduced computational cost.